# A Reanalysis of the FDA’s Benefit-Risk Assessment of Moderna’s mRNA-1273 COVID Vaccine: For 18–25-Year-Old Males, Risks Exceeded Benefits Relative to Hospitalizations

### v1

@@ 5. Validating the Model
In this section, we discuss our model-validation work of Supplement 2, where we test our model as well as the FDA’s using COVID-hospitalizations data from Ontario over the evaluation period 1/1/22 – 5/31/22 and show that our model produces projections of the hospitalizations-prevented benefit of vaccination that are much more accurate than those of the FDA [9]. We test our model in two ways: (i) showing that it fits Ontario COVID-hospitalization data with inputs at plausible levels for Ontario over the evaluation period (see Section 8 of Supplement 2), and (ii) using Ontario COVID-hospitalization data to predict corresponding data for the U.S. and then comparing those predictions to ours (see Section 7 of Supplement 2). We validate our model using Ontario data for many reasons; e.g., unlike the CDC, Public Health Ontario chose to publish incidental COVID-hospitalization rates (starting 1/8/2022). We discuss other reasons in Section 3 of Supplement 2.
Because Public Health Ontario does not provide sex-specific COVID-hospitalizations data, to test our model using its data, we compute model outputs assuming no sex-related differences in COVID infection and hospitalization risk, so that $H_{r}=0.0022$, and, for Scenario D, $F_{pi}=0.76$ (see Supplement 1, Sections 2 and 3). If $H_{r}$ and $F_{pi}$ are increased to $0.0022$ and $0.76$ in Scenario D, leaving all other variables unchanged, then 231 is the hospitalizations-prevented benefit per million vaccinations.
Based on COVID hospitalization counts provided by Public Health Ontario as well as corresponding incidental hospitalization rates, we estimate that vaccinating one million 18–25-year-old Ontarians prevented 152 COVID hospitalizations in Ontario over the evaluation period—see Table S2.5. Because seroprevalence studies suggest that over the evaluation period, the COVID-infection rate in Ontario exceeded that in the U.S., we would anticipate a hospitalizations-prevented benefit of less than 152 per million in the U.S. provided we assume that 18–25 year-olds in the U.S. and in Ontario have equivalent comorbidity profiles. Making adjustments for better health of Ontarians, ignoring the greater infection-level rise in Ontario, and assuming all the vaccinated were 14 days postdose 2 on 1/1/22, we estimate that at most 214 evaluation-period hospitalizations were prevented by one million COVID vaccinations of 18–25 year-olds in the U.S. [Supplement 2, Table S2.7]. [^23] In addition to the greater infection-level rise in Ontario, other reasons why we view the rate of 214 likely to be an upper bound on the actual rate include the following: our ignoring evidence that incidental COVID hospitalization rates are higher among the young, ignoring evidence that hospitalization rates are higher among females, and applying a better-health adjustment such that “the better health of Ontarians [relative to Americans] is more likely overstated than understated”—see Section 4 of Supplement 2. Observe that our most likely scenario’s predicted benefit of 231 hospitalizations prevented per million vaccinations (assuming equivalent IHR for males and females) is only about 1.08 times larger than the estimated actual benefit of 214 [Supplement 2, Table S2.7], while FDA’s most likely scenario’s predicted benefit of 4,766 hospitalizations prevented per million vaccinations is over 22 times larger. Moreover, had the FDA assumed equivalent infection and hospitalization rates among males and females, then its corresponding benefit projection likely would have been over 6,000 hospitalizations prevented per million vaccinations.
@@
Ontario data also supports the bounding analysis of Scenario I from Section 4.1 in which 155.82 in 1 million is used as an upper bound on the evaluation-period COVID-hospitalization risk for a typical unvaccinated male Bob having prior COVID infection and no comorbidities. Ontario data suggests that 151 in 1 million is an overestimate of Bob’s evaluation-period COVID hospitalization risk—see Section 2.3 of Supplement 2.
@@ 6. Discussion
Our reanalysis of the FDA’s benefit-risk assessment [9] of Moderna mRNA-1273 primary-series vaccination reveals substantial shortcomings in the modeling that supported the agency’s finding “the benefits of vaccination with the Moderna COVID-19 vaccine clearly outweigh the risks” even for males of ages 18–25 years. By accounting for prior-infection protection and using evidence-based model inputs, we find in our Scenarios B–E that risks outweighed benefits for 18–25-year-old males, assuming hospitalizations are the principal endpoint of analysis. In particular, for our scenarios predicting plausible infection-level increases over the evaluation period (Scenarios B, D, and E), our model suggests that vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myo/pericarditis (VAM/P) alone compared to COVID hospitalizations prevented. [^24] The range 16%–63% is consistent with our model-validation work discussed in Section 5: recall we estimate that at most 214 evaluation-period hospitalizations were prevented by one million mRNA-1273 vaccinations of 18–25-year-old males in the U.S. [Supplement 2, Table S2.7], and note that our corresponding VAM/P hospitalization projection of 268 is a little over 25% higher than 214. Moreover, for the subpopulation of 18–25-year-old males having prior-infection protection and no comorbidities, we have established via Scenarios I and II of Section 4.1 that the choice to mRNA-1273 vaccinate created a greater risk of hospitalization (owing to VAM/P or an evaluation-period breakthrough infection) than the risk of COVD hospitalization if unvaccinated.
The shortcomings of the FDA’s analysis included not accounting for prior-infection protection as well as incidental hospitalizations, and overestimating the COVID-hospitalization rate for 18–25-year-old males by assuming it to be equivalent to the rate for 18–45-year-old males. On the risk side, as discussed in Section 2.3, the FDA relied on a very small internal dataset (21 second-dose events for mRNA-1273 [44]) to determine mRNA-1273 VAM/P risk rather than relying on findings from regulatory agencies from 5 other countries, from the CDC’s VSD system, from major international studies [42, 43], and from a U.S. study [45]. These shortcomings contributed to the FDA’s obtaining in its most likely scenario a benefit-risk ratio of 43.33 for mRNA-1273 vaccination of 18–25-year-old males, a ratio that is over 60 times more favorable than that of our most likely scenario, 0.67.
6.1. The strength of prior-infection protection
The FDA decided to not account for prior-infection protection: “We did not factor in protection from potential exposure or infection with COVID-19 among unvaccinated individuals into our analysis since the degree of this protection and its time-dependency were not well understood at the time of the analysis” [9, Section 2.2.1.1]. However, the protection provided by mRNA-1273 vaccination against the Omicron variant and its time-dependency were also not well understood at the time of analysis (3rd week of January 2022). Nevertheless, the FDA assumed in every one of its five Omicron-based benefit-risk scenarios that mRNA-1273 vaccination would provide for at least 5 months’ protection against cases with 30% effectiveness and against hospitalization with 72% effectiveness. Even for its Delta-based scenario of [9] the FDA assumed prior-infection would provide no protection, despite (i) the CDC study [18] indicating that prior infection provides superior protection against Delta than does mRNA vaccination and (ii) a CDC Science Brief stating, “Available evidence shows that fully vaccinated individuals and those previously infected with SARS-CoV-2 each have a low risk of subsequent infection for at least 6 months” ([17], 10/29/21).
@@ 6.1. The strength of prior-infection protection
As noted in Section 2.2, the FDA’s assumption that prior infection would provide no protection against Omicron infection and hospitalization is contrary to immunological expectations and studies (e.g., [16, 17, 18, 27, 28]) available to the FDA at the time of its mRNA-1273 analysis all suggesting that prior-infection protection would likely equal or exceed vaccine protection among the COVID-naïve. As noted in Section 3.3, we have assumed that protection against hospitalization provided by vaccination is 77% while that provided by prior infection is 88.9%. A systematic review [25] involving 21 vaccine-effectiveness studies from the period 3 December 2021 – 7 April 2022 reports,
@@
In contrast, the meta-analysis [22] estimates prior-infection protection against severe Omicron to be over 80% through 6 months after infection, while the meta-analysis [23, Table S2] estimates protection to be over 88% through 45 weeks after infection, with the former analysis based on studies from the period 1/1/22 – 6/1/22 and the latter, on studies published up to 9/31/22. In early January of 2022, the minimum reasonable assumption the FDA might have made is that prior-infection protection equals vaccine protection (as in our Scenario E); instead, in all Omicron-based scenarios of [9], the agency ascribed to vaccination 72% effective protection against hospitalization while ascribing 0% effective protection to prior infection.
The FDA’s reliance on the UK Health Security Agency for its estimates of vaccine effectiveness against Omicron [34], instead of relying on the CDC, is surprising. The CDC might have had precisely the data the FDA needed for its mRNA-1273 assessment. For example, if the CDC extended its May–November 2021 study [18] through the end of December 2021, then the CDC might have had estimates of vaccine as well as prior-infection effectiveness against Omicron at the time of the FDA’s mRNA-1273 analysis. The FDA’s assessments of BNT162b2 [8] and mRNA-1273 [9] suggest that the FDA does not necessarily request from the CDC data needed to conduct rigorous analyses. For both assessments, the FDA failed to obtain data on incidental COVID hospitalizations. For both assessments, the FDA claims that age/sex specific VAM/P information in the CDC’s VSD system is not available; e.g., Funk et al. claim “[W]e have no access to VSD data for smaller age/sex subgroups” [8, Section 2.3.2.1]. Finally, for its mRNA-1273 assessment, the FDA failed to request that the CDC disaggregate COVID hospitalization data for the age range 18–45 into data for the subranges 18–25, 26–35, and 35–45. Also troubling is the withholding of and underreporting of ICU-admissions data for VAM/P in, respectively, [8] and [9] (see Section 2.4) as well as willingness to use international data [34] for the benefit analysis of [9] while eschewing international data relating to risks (e.g., VAM/P data from Public Health Ontario [3, 42] and postvaccination myocarditis data from England ([6, 43]).
6.2. Summary
We have shown that extending the FDA’s model to account for prior-infection protection as well as incidental COVID hospitalizations and making evidence-based model-input assumptions is sufficient to invert the calculus of seemingly overwhelming benefit of mRNA-1273 vaccination towards excessive risk of VAM/P in the general population of 18–25-year-old males. The failure of the FDA author team of [9] to perform a rigorous benefit-risk analysis led to COVID-vaccination recommendations that both failed reflect “The health condition [and prior-infection status] of individuals may remarkably impact the B-R profile and its evaluation” ([9, Section 4]) and that supported vaccination mandates enforced by many institutions placing certain subgroups at a greater risk of hospitalization for a serious adverse event owing to COVID vaccination than for hospitalization owing to COVID infection—while on an absolute basis this risk difference is small, institutional responsibility for adverse events owing to vaccination is not.
7. Recommendations
@@
Our modeling provides a call to action to compel our policy-forming health agencies to employ multi-parameter models that can better stratify vaccination recommendations. To protect vaccine acceptance and promote public trust, [^25] A survey from the Robert Wood Johnson Foundation and Harvard School of Public Health released May of 2021 found only 37% of respondents had a great deal of trust in the FDA, while 38% somewhat trusted the agency, and 24% described their level of trust “Not very much/Not at all” [74]. the FDA must adopt a rigorous framework for benefit-risk assessment of new vaccines and revisit recommendations as new data becomes available. In particular, the framework must incorporate consideration of the protection that prior infection affords against development of severe disease.
@@
A re-alignment of priorities of the public-health agencies of the United States is necessary. Two years after the first cases SARS-CoV-2 were identified in the U.S., the FDA decided to “not investigate the benefits and risks [of mRNA-1273 vaccination] for specific subpopulations, such as those with comorbidities, due to limited information for these populations” [9, Section 4]. Also, public-health agencies should have prioritized understanding the effectiveness of prior-infection protection against reinfection and hospitalization. That said, in Pfizer’s phase-3 trial for BNT162b2, 567 participants in the placebo group and 526 in the treatment arm had evidence of prior COVID-19 infection. In each of these cohorts, there was only 1 reinfection (or ¡0.2% for both) according to the primary endpoint definition [75, p. 27]. This alone should have prompted health agencies to aggressively pursue further data on protection from prior infection to stratify risk and recommendations. We note that by late fall of 2022, the percentage of the U.S. population having experienced at least one COVID infection was estimated to be 95% by an IHME study [76] and 94% by a Harvard study [77].
Given that nearly all persons in the United States now have significant protection from severe COVID conferred by prior-infection, there is an urgent need to reassess whether the following current blanket CDC recommendations [78] for COVID-19 vaccination advance public health: “ Everyone ages 6 months and older should get a 2024–2025 COVID-19 vaccine.” Compare the CDC’s policy to that of the UK: “All those aged 65 or over, all those in care homes for older people, and those aged 6 months or over who are in a clinical risk group will be eligible for vaccination this autumn” [79].
The urgency of the need for reassessment is reinforced by the fact that Pfizer’s and Moderna’s reformulated COVID-19 vaccines are available only via emergency-use authorization for individuals 6 months through 11 years of age (despite the expiration of the COVID-19 public health emergency 11 May 2023).
8. Limitations
@@
We have assumed that young males are hospitalized for COVID at a higher rate than young females, as did the FDA in its benefit-risk assessments [8] and [9]. Data from [8] and [54] suggests this rate difference is substantial; however, the rates of [8] and [54] include incidental hospitalizations and incidental rates for males may differ from those for females. We have assumed, as did the FDA, that protection provided by vaccination and prior infection against hospitalization and infection is essentially equivalent for males and females, and this may not be the case. Also, we have based our assessment of the difference in hospitalization rates between young males and females on data in Table 2 of [8] representing mostly pre-Delta cases. Relative to Omicron, sex-based differences in hospitalization rates might have changed. However, the rate differences provided by Table 1 of [54], based on two years of data including that of the initial Omicron wave, are consistent with those of Table 2 of [8].
@@
COVID-hospitalizations data from Public Health Ontario (see Supplement 2) and from the State of Connecticut (see the appendix of Supplement 1) suggests that that our Omicron IHR value $0.0014$ for 18–25-year-old males is a reasonable estimate for the Omicron IHR for all 18–25 year-olds (general population mix of males and females). Thus, even if we assume males and females in this age range are hospitalized for COVID at roughly equivalent rates, our IHR value of 0.14% for 18–25-year-old males appears to be a good estimate.
@@
As noted in Section 3.3, based on data available before 1/22/22, we arrived at an Omicron IHR estimate for all 18–25 year-olds of 0.22%, not 0.14%. One possible explanation is that we may have underestimated the incidental COVID-hospitalization rate for 18–25 year-olds. Our IHR estimate for 18–25 year-olds of 0.22% results from assuming an incidental COVID-hospitalization rate of 40%, which is the rate suggested by [29] and [31] for hospitalizations with COVID among patients of all ages. However, there is strong evidence that incidental COVID-hospitalization rates are higher among the young than among the elderly; see, e.g., [30], [80], and Supplement 1. Figure 2 of [30] suggests an incidental rate of over 60% among 0–18 year-olds from April of 2020 through May of 2021. COVID-hospitalization data from the CDC and the state of Connecticut, discussed in the appendix of Supplement 1, suggests an incidental rate of about 72% among 15–24 year-olds in Connecticut during calendar year 2020. Assuming an incidental rate of about 62.2% (instead of 40%) transforms our IHR estimate of 0.22% into the estimate 0.14%.
In terms of applying our model to an emerging pathogen and vaccine campaign in the future, our functions $U$ and $V$ require reasonably accurate inputs to produce reasonably accurate projected hospitalizations as outputs. In particular, if public-health agencies do not provide accurate rates of hospitalizations (requiring treatment for the infection) based on age, vaccination-status, sex, and prior-infection status, then obtaining accurate infection-hospitalization rates and rate-reduction factors becomes difficult.
In applying our improvement of the FDA’s model for the general population of 18–25-year-old males, we constrained ourselves to using only information available at the time of the FDA’s benefit-risk analysis of mRNA-1273 .
@@
Our assessment of risks associated with vaccination, like the FDA’s, ignores known rare but potentially severe vaccine-related adverse events other than myo/pericarditis that may require hospitalization, for example, anaphylaxis [58], Multisystem Inflammatory Syndrome [59, 60], Guillain Barré Syndrome [81], posterior orthostatic tachycardia syndrome (POTS) [82], retinal vascular occlusion [83], and cerebral venous sinus thrombosis (CVST) [64]. Our risk assessment as well as the FDA’s ignores the possibility that COVID-19 vaccination has long-term health effects; for instance, the FDA’s package-insert information [84, Section 13] for Moderna’s mRNA-1273 begins, “SPIKEVAX has not been evaluated for carcinogenic, mutagenic potential, or impairment of male fertility in animals.”
Finally, we have carried out a retrospective analysis illustrating how the FDA might have conducted its mRNA-1273 benefit-risk assessment in a more rigorous way. Data we rely on for assessing benefit and risks of vaccination pertains to the original, monovalent formulation of Moderna’s COVID vaccine, having a two-dose vaccination schedule.
@@
Acknowledgments: We thank Retsef Levi for providing comments that improved both the structure and content of this paper.
Links to Supplements:

### v3

@@ 4.3. Validating the Model
In Supplement 2, we use Ontario COVID-hospitalizations data from the evaluation period 1/1/22– 5/31/22 to show that our model’s projections of Table 6 for the hospitalizations-prevented benefit of vaccination for 18–25-year-old males are much more accurate than those of the FDA [11]. [^14] We use Ontario data for many reasons; e.g., unlike the CDC, Public Health Ontario chose to publish incidental COVID-hospitalization rates (starting 1/8/2022) [30]; see [Supplement 2, Section 3] for other reasons. We test our model in two ways: (i) showing that it fits Ontario COVID-hospitalization data with inputs at plausible levels for Ontario over the evaluation period [Supplement 2 Section 8], and (ii) using Ontario COVID-hospitalization data to predict corresponding data for the U.S. [Supplement 2, Section 7]. For (i), we fit our model to Ontario data for the general-population mix of 18–25-year-old male-and-female Ontarians by using Mathematica to solve a sequence of least-squares problems depending on the value of $H_{r}$. For $H_{r}=0.0011$ and $0.0010$, we obtain model inputs in plausible ranges for Ontario that yield a perfect modeling fit and note that these values of $H_{r}$ are consistent with COVID hospitalization data from the state of Connecticut. A brief discussion of (ii) is provided in the next paragraph.
Based on COVID hospitalization counts from Public Health Ontario as well as corresponding incidental hospitalization rates, we estimate that vaccinating one million 18–25-year-old Ontarians prevented 152 COVID hospitalizations over the evaluation period—see Table S2.5. Making various adjustments e.g., accounting for the better health of Ontarians vs. Americans, we obtain an upper-bound estimate that 214 evaluation-period hospitalizations were prevented by one million full COVID vaccinations of 18–25–year-old males in the U.S.—see [Supplement 2, Table S2.7] and the discussion that follows it. Observe that 214 is an upper bound for our projections of Scenarios B–E (see Table 6) but not an upper bound for any of the FDA’s projections.
@@
Ontario data also supports the definitiveness of the results of Table 7 above given the male has no comorbidities, suggesting that 151 in 1 million is an overestimate of his evaluation-period COVID hospitalization risk [Supplement 2, Section 2.3].
@@ 5. Discussion
By accounting for prior-infection protection and using evidence-based model inputs, we find in our Scenarios B–E of Table 6 that risks of mRNA-1273 vaccination outweighed benefits for the general population of 18–25-year-old males, assuming hospitalizations are the principal endpoint of analysis. In particular, for our scenarios predicting plausible infection-level increases over the evaluation period (Scenarios B, D, and E), our model suggests that vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myo/pericarditis (VAM/P) alone compared to COVID hospitalizations prevented. [^15] The range 16%–63% is consistent with our model-validation work discussed in Section 4.3: recall we estimate that at most 214 evaluation-period hospitalizations were prevented by one million mRNA-1273 vaccinations of 18–25-year-old males in the U.S. [Supplement 2, Table S2.7], and note that our corresponding VAM/P hospitalization projection of 268 is a little over 25% higher than 214. Recalling that the FDA found, “[T]he benefits of vaccination with the Moderna COVID-19 vaccine clearly outweigh the risks, even among the male adolescent population” [11, Section 4], we have demonstrated that the conclusions of a benefit-risk assessment for vaccination may be dramatically impacted by accounting for the benefits conferred by prior infection. We have also demonstrated in Tables 79 how the benefit-risk picture for vaccination may vary based on an individual’s prior-infection and comorbidity status.
Consistent with our model’s finding that for 18–25-year-old males in the U.S. the risks of mRNA-1273 vaccination outweighed the benefits relative to hospitalizations, we note that France suspended mRNA-1273 use on 10/15/2021, making it available again on 11/8/21 but at half-dose and only for those over 30 years [73, p. 4].
We emphasize that the cardiac risk of COVID infection is incorporated into our modeling because our estimates of COVID hospitalizations include those for COVID-related cardiac issues. We address here assertions such as “Importantly, the risk of myocarditis after any dose of COVID-19 vaccine, in any population, is substantially lower than the risk of myocarditis after SARS-CoV-2 infection,” which appeared, without supporting references, in an April 2024 editorial of the European Heart Journal [74]. Contrary to the editorial’s assertion, Patone et al. find that “In men younger than 40 years old, the number of excess myocarditis events [leading to hospitalization] per million people was higher after a second dose of mRNA-1273 than after a positive SARS-CoV-2 test (97 [95% CI, 91–99] versus 16 [95% CI, 12–18])” [6]. Karlstad et al. also find significantly higher risk of hospitalized myocarditis after dose 2: 183.9 per million 2nd mRNA-1273 doses [5, Table 2] vs. 13.7 per million positive tests [5, eTable 7] among males 16–24. Moreover, these myocarditis incidence rates expressed in cases per million positive tests overstate the risk of myocarditis after COVID infection expressed in cases per million COVID infections (because many infections are not confirmed by positive tests, but essentially all infections among  those seeking care for myocarditis are confirmed by positive tests) [75].  Consistent with [75], the authors of [76] note, “[T]he attributable risk estimates for COVID-19 [associated myocarditis] used laboratory confirmed cases as the denominator and will be affected by the proportion of all SARS-CoV-2 infections captured by testing, precluding a direct comparison with vaccine-associated attributable risks.”
The strength of prior-infection protection
@@
As noted in Section 2.2, the FDA’s assumption that prior infection would provide no protection against Omicron is contrary to immunological expectations and studies (e.g., [15, 16, 17, 19, 26, 27]) available at the time of the FDA’s mRNA-1273 analysis suggesting that prior-infection protection would likely equal or exceed vaccine protection among the COVID-naïve. As noted in Table 5 footnotes, we have assumed that vaccination’s effectiveness against hospitalization is 77% while that provided by prior infection is 88.9%. A systematic review [24] involving 21 vaccine-effectiveness studies from the period 12/3/2021 – 4/7/2022 reports,
@@
In contrast, the meta-analysis [21] estimates prior-infection’s effectiveness against severe Omicron to be over 80% through 6 months after infection, while the meta-analysis [22, Table S2] estimates its effectiveness to be over 88% through 45 weeks after infection, with the former estimate based on studies from 1/1/22–6/1/22 and the latter, on studies published by 9/31/22. In mid-January of 2022, the minimum reasonable assumption the FDA might have made is that prior-infection protection would equal vaccine protection (as in our Scenario E); instead, in all Omicron-based scenarios of [11], the agency ascribed to vaccination 72% effective protection against hospitalization while assuming prior infection offers no protection.
5.1. Recommendations
@@
Our modeling provides a call to action to policy-forming health agencies to employ multi-variable models like ours that can better stratify vaccination recommendations. To protect vaccine acceptance and promote public trust, the FDA must adopt a rigorous framework for benefit-risk assessment of new vaccines and revisit recommendations as new data becomes available. In particular, the framework must incorporate consideration of the protection that prior infection affords against development of severe disease.
@@
We also recommend United States public-health agencies reassess their priorities. Two years after the first cases of COVID were identified in the U.S., the FDA decided to “not investigate the benefits and risks [of mRNA-1273 vaccination] for specific subpopulations, such as those with comorbidities, due to limited information for these populations” [11, Section 4]. Also, public-health agencies had reason to prioritize understanding the effectiveness of prior-infection protection and had evidence that it was likely to be high. For example, in Pfizer’s phase-3 trial for BNT162b2, 567 participants in the placebo group and 526 in the treatment arm had evidence of prior COVID-19 infection. In each of these cohorts, there was only 1 reinfection (or <0.2<0.2% for both) [77, p. 27].
5.2. Limitations
@@
Our principal aim has been to illustrate how prior-infection protection, as well as comorbidity status, can be integrated into benefit-risk modeling of a two-dose vaccination series. In applying our model to extend and improve the FDA’s mRNA-1273 assessment [11] relative to its hospitalizations endpoint, we have assumed that young males are hospitalized for COVID at a lower rate than young females, as did the FDA in [10] and [11]. Data from [10] and [54] suggests this rate difference is substantial; however, the rates of [10] and [54] include incidental hospitalizations and incidental rates for males may differ from those for females. We have assumed, as did the FDA, that vaccination and prior infection offer essentially equivalent protection for males and females. Also, we have based our assessment of the difference in hospitalization rates between young males and females on data in Table 2 of [10] representing mostly pre-Delta cases. Relative to Omicron, sex-based differences in hospitalization rates might have changed. However, the rate differences provided by Table 1 of [54], based on two years of data including that of the initial Omicron wave, are consistent with those of Table 2 of [10].
@@
COVID-hospitalizations data from Ontario [Supplement 2] and from Connecticut [Supplement 1, Appendix 1] suggests that our Omicron IHR value $0.0014$ for 18–25-year-old males is a reasonable estimate for the Omicron IHR for all 18–25 year-olds (general population mix of males and females). Thus, even if we assume males and females in this age range are hospitalized for COVID at roughly equivalent rates, our IHR value of 0.14% for 18–25-year-old males appears to be a good estimate.
@@
In Section 2 of Supplement 1, based on data available before 1/22/22, we obtain an Omicron IHR estimate for 18–25 year-olds (general-population mix of males and females) of 0.22%, not 0.14%. One possible explanation is that we may have underestimated the incidental COVID-hospitalization rate for 18–25 year-olds. Our IHR estimate for 18–25 year-olds of 0.22% results from assuming an incidental Omicron-hospitalization rate of 40%, a rate suggested by [28] and [30] for hospitalizations with COVID among patients of all ages. However, as noted in Section 2.2 (and discussed in detail in Supplement 1), there is strong evidence that incidental COVID-hospitalization rates are higher among the young than among the elderly. In particular, COVID-hospitalization data discussed in Appendix 1 of Supplement 1 suggests an incidental rate of over 70% among 15–24 year-olds in Connecticut during calendar year 2020. Assuming an incidental rate of about 62.2% (instead of 40%) transforms our IHR estimate of 0.22% into the estimate 0.14%.
For a vaccine having high effectiveness against infection that does not quickly wane, vaccination can have an important benefit not captured by our model—providing an enduring reduction in the level of infection in the population that may allow some members to avoid infection entirely through the contribution of vaccination to herd immunity.
@@
Our risk assessment, like the FDA’s, ignores known, rare vaccine-related adverse events other than myo/pericarditis that may require hospitalization. Finally, our risk assessment, like the FDA’s, ignores the possibility that COVID-19 vaccination may have long-term health effects; for instance, the FDA’s package-insert information for Moderna’s mRNA-1273 begins, “SPIKEVAX has not been evaluated for carcinogenic, mutagenic potential, or impairment of male fertility in animals” [78, Section 13].
@@
Acknowledgment: We thank Retsef Levi for providing comments that improved both the structure and content of this paper.
Supplements: Supplement 1, Supplement 2